\begin{thebibliography}{99}

%\selectlanguage{russian}

\bibitem[Be88]{Be88} A. Bendukidze, Fermat looks for extrema (in Russian), Kvant, 1988, N10.
%Бендукидзе А., Ферма ищет экстремумы, Квант, 1988, N10.
%\url{http://kvant.ras.ru/1988/10/ferma_ishchet_ekstremumy.htm}

\bibitem[DFMS]{DFMS} S. Doledenok, M. Fadin, A. Menshchikov, A. Semchankau, The $pqr$-method,
\linebreak
\url{http://www.turgor.ru/lktg/2016/3/index.htm}.

\bibitem[FT]{FT} D. Fuchs, S. Tabachnikov,  Mathematical Omnibus. AMS, 2007.
\linebreak
\url{http://www.math.psu.edu/tabachni/Books/taba.pdf}.

\bibitem[Go10]{Go10} M. Gorelov, Why graphs are useful (in Russian), Kvant, 2010, N3, 44--~47.
%Горелов М., О пользе графиков. Квант,

\bibitem[Ta88]{Ta88} S. Tabachnikov, Geometry of equations (in Russian), Kvant, 1988, N10.
%Табачников С., Геометрия уравнений. Квант,

\bibitem[Po84]{Po84} L. Pontryagin, Cubic parabola (in Russian) Kvant, 1984, N3.
%Понтрягин Л., Кубическая парабола. Квант

\bibitem[ZSS]{ZSS}
Elements of Mathematics As a Sequence of Problems (in Russian), editors: A. Zaslavsky, A. Skopenkov, M. Skopenkov, Moscow, MCCME, 2018.
\linebreak
\url{http://www.mccme.ru/circles/oim/materials/sturm.pdf}

%\bibitem[Pr07]{Pr07}
%В.В. Прасолов, Задачи по алгебре, арифметике и анализу, М.: МЦНМО, 2007.
%\linebreak
%\url{http://www.mccme.ru/prasolov}

%\bibitem[K]{K} V. Kolosov, Theorems and Problems of Algebra, Combinatorics and Number Theory (in Russian), Gelios, Moscow, 2001.
%\bibitem[PS]{PS} V. Prasolov and Yu. Solovyov, Elliptic Functions and Elliptic Integrals, AMS, 1997.
%Russian version: \url{www.mccme.ru/~prasolov}.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage[russian,english]{babel}
\usepackage{amsfonts,amsmath,amssymb,amscd,amsthm,url}
\long\def\comment#1\endcomment{}

\def\sgn{\mathop{\fam0 sgn}}

    \theoremstyle{theorem}
         \newtheorem{Theorem}{Theorem}
         \newtheorem{Lemma}[Theorem]{Lemma}

\newtheoremstyle{mydefinition}
  {3pt}
  {3pt}
  {\normalfont}
  {\parindent}
  {\bfseries}
  {.}
  { }
  {}

\theoremstyle{mydefinition}

\begin{document}


\selectlanguage{english}

\centerline{\uppercase{\bf How Fermat found extrema}
\footnote{I am grateful to A. Doledenok, M. Skopenkov and A. Sgibnev for useful discussions.}
}

\bigskip
\centerline{\bf A. Skopenkov \footnote{\url{www.mccme.ru/~skopenko}.
Moscow Institute of Physics and Technology and Independent University of Moscow.
Supported in part by the D. Zimin Dynasty Foundation and Simons-IUM fellowship.
}}

\bigskip
{\bf Abstract.}
We present a short elementary proof of the well-known criterion for a cubic polynomial to have three real roots.
The proof is based on Fermat's approach to calculus for polynomials. 
This approach illustrates the idea of a derivative rigorously but without technical $\varepsilon$-$\delta$ language.
The note is accessible to high-school students.



\bigskip
In elementary school one learns the following.

{\it Let $a$ and $b$ be real numbers.
Then the following conditions are equivalent:

$\bullet$ there are real numbers $x,y$ such that $a=x+y$ and $b=xy$;

$\bullet$ the equation $t^2-at+b=0$ has a real root;

$\bullet$ $4b-a^2\le0.$}

In this note we illustrate the idea of a derivative by proving a well-known generalization to {\it three} numbers
(which might be proved by Fermat or earlier).

\begin{Theorem}\label{t:fermat}
Let $a,b,c$ be real numbers.
There are real numbers $x,y,z$ such that
$$a=x+y+z,\quad b=xy+yz+zx\quad\text{and}\quad c=xyz$$
if and only if
$$4\left(b-\dfrac{a^2}3\right)^3+27\left(c-\dfrac{ab}3+\dfrac{2a^3}{27}\right)^2\le0.$$
\end{Theorem}

This result is very useful; for recent applications to elementary inequalities see \cite{DFMS} and references therein.
Not only this result, but simple exposition of its proof is hopefully interesting.
We do not use the notion of a derivative.
However, we illustrate the idea of a derivative rigorously but without technical $\varepsilon$-$\delta$ language, cf. \cite[\S8.1, \S8.2]{ZSS}.
So this exposition might be useful for introductory courses on analysis.
Unfortunately, this exposition is not so well-known, cf. \cite{Po84}.\footnote{E.g. in June 2016 it was suggested to add this proof to \cite{DFMS} in order to make Theorem \ref{t:fermat}
and its applications more accessible.
However, the authors found this proof `too complicated'. 
The first version of this note (\url{https://arxiv.org/pdf/1610.05968v1.pdf}) is rejected from posting on \url{http://www.turgor.ru/lktg/2016} by the authors of \cite{DFMS} and S. Dorichenko.}
For development of the `graphs of functions' idea see \cite{FT, Go10, Ta88} (in particular, one can apply this idea to the $pqr$-lemmas of \cite{DFMS}).

There is an alternative proof of Theorem \ref{t:fermat} using complex numbers and calculation of the {\it discriminant} of cubic polynomial in terms of coefficients, see e.g. \cite[\S2 and solutions of problems 6-22]{DFMS}.
Although that proof is longer, it illustrates another interesting important ideas.
In \cite{Ta88} a geometric interpretation (but not a proof) of Theorem \ref{t:fermat} is presented.

\medskip
{\it A reformulation of Theorem \ref{t:fermat} in terms of cubic polynomials.}
By Vieta Theorem (i.e. by direct calculation) the condition on $x,y,z$ is equivalent to the following equality of polynomials:
$$t^3-at^2+bt-c=(t-x)(t-y)(t-z).$$

{\it A reduction of Theorem \ref{t:fermat} to the case $a=0$.}
Take $u:=t-\frac a3$. Then $t=u+\frac a3$, so
$$t^3-at^2+bt-c=u^3+\left(b-\dfrac{a^2}3\right)u-\left(c-\dfrac{ab}3+\dfrac{2a^3}{27}\right).$$
Thus it suffices to prove Theorem \ref{t:fermat} for $a=0$.

\begin{proof}[Proof of Theorem \ref{t:fermat} for $a=0$, $b\ge0$]
Assume that the required $x,y,z$ exist.
Since $b\ge0$, the function $t^3+bt-c$ is strictly increasing as a sum of increasing functions, one of them strictly increasing.
Then the equation $t^3+bt-c=0$ has at most one real root.
This and $a=0$ imply that $x=y=z=0$, hence $4b^3+27c^2=0$.

If $4b^3+27c^2\le0$, then $b=c=0$, so we can take $x=y=z=0$.
\end{proof}

{\it Heuristic considerations for investigation of the function $f(t):=t^3-3t$} (formally they are not used in the proof).
The function $f$ is called {\it strictly increasing} on an interval if $f(t_1)>f(t_2)$ for each different $t_1>t_2$ from the interval.
This is equivalent to $\varphi(t_1,t_2)>0$ for each different $t_1,t_2$ from the interval, where
$$(*)\qquad\varphi(t_1,t_2):=\dfrac{f(t_1)-f(t_2)}{t_1-t_2}=t_1^2+t_1t_2+t_2^2-3.$$
If these conditions hold for each two `close' $t_1,t_2$, then by transitivity they hold for each two $t_1,t_2$.
So we guess that the `boundary points' of the intervals on which $f$ is monotonous are the roots of the equation $t^2+tt+t^2-3=0$.
These roots are $\pm1$.
(This is analogous to \cite{Be88}; considering a simple example before general method makes the method more accessible.)

\begin{Lemma}\label{l:incr}
The function $f(t):=t^3-3t$ is strictly increasing on $(-\infty,-1]$, is strictly decreasing on $[-1,1]$ and is strictly increasing on $[1,+\infty)$.
\end{Lemma}

\begin{proof} Denote $\varphi(t_1,t_2)$ by the formula (*).
Then $\varphi(t_1,t_2)>0$ for each different $t_1,t_2\ge1$.
Hence $f(t)$ strictly increases on $[1,+\infty)$.
Analogously the other two statements are proved.
\end{proof}

\begin{proof}[Proof of Theorem \ref{t:fermat} for $a=0$, $b<0$]
Taking
$u:=t\sqrt{-b/3}$ we may assume that $b=-3$.

First assume that the required $x,y,z$ exist.
Since $-3=b<0=a$, the case $x=y=z$ is impossible, so the equation $f(t)=t^3-3t-c=0$ has at least two real roots.

Denote $f(t):=t^3-3t-c=t(t^2-3)-c$.
Then $t_+:=1+\max\{3,|c|\}>1$ and $f(t_+)>0$.
Analogously, there is $t_-<-1$ such that $f(t_-)<0$.
So by Lemma \ref{l:incr} and the Intermediate Value Theorem, the equation $f(t)=0$ has at least two real roots if and only if $f(-1)$ and $f(1)$ `have different signs', i.e. $f(-1)f(1)\le0$.
We have $-f(t)=c-t(t^2-3)$, so
$$f(-1)f(1)=(c+2\cdot1)(c-2\cdot1)=c^2-4=(4b^3+27c^2)/27.$$

Now assume that $4b^3+27c^2=27(c^2-4)\le0$.
If $c^2<4$, the above paragraph shows that the equation $f(t)=0$ has three real roots, and we take them as $x,y,z$.
If $c^2=4$, we take $x=y=-\sgn c$ and $z=2\sgn c$ (it  should be clear how to guess these formulas).
\end{proof}

{\bf Exercise.} Find maximal intervals on which the following function is strictly increasing (decreasing):

(a) $f(t)=t^4-4t$. \quad (b) $f(t)=t^4-12t^3+22t^2-24t+10$.


\end{document}